Ram13:

Find the four angles that define the fourth roots of the number z1 = 1 + sqrt 3 * i

surjithayer:

\[z1 =r (\cos \theta+\iota \sin \theta) \] \[r \cos \theta=1\] \[r \sin \theta=\sqrt {3}\] square and add \[r^2=1+3=4\] r=2 divide \[\tan \theta=\sqrt{3}=\tan 60=\tan (180n+60)\] \[\theta=180n+60\] \[z1=r(\cos \theta+\iota \sin \theta)=re^{\iota \theta}\] \[z1=2e^{\iota (180n+60)}\] \[z1^{\frac{ 1 }{ 4 }}=2^{\frac{ 1 }{ 4}}e ^{\frac{ 180n+60 }{ 4 }}\] \[z1^{\frac{ 1 }{ 4 }}=2^{\frac{ 1 }{ 4 }}e ^{45n+15}\] put n=0,1,2,3 fourth roots are \[2^{\frac{ 1 }{ 4 }}e ^{15 \iota },2^{\frac{ 1 }{ 4 }}e ^{60 \iota} ,2^{\frac{ 1 }{ 4}}e ^{150 \iota},2^{\frac{ 1 }{ 4 }}e ^{195 \iota }\]

surjithayer:

correction last line fourth roots are \[2^{\frac{ 1 }{ 4 }}e ^{15 \iota},2^{\frac{ 1 }{ 4 }}e ^{ 60 \iota},2^{\frac{ 1 }{4 }}e ^{105 \iota },2^{\frac{ 1 }{ 4 }}e ^{150 \iota }\]